# Framing Unionization on Facebook: Communication around Representation Elections in the United States

### v2

@@ Introduction
https://github.com/ariannap13/laborunions. Models are available on HuggingFace as ariannap22/roberta˙finetuned˙fb˙unions and ariannap22/roberta˙finetuned˙multilabel˙unions
@@
  • $\text{{pre}}_{fi}$: mean proportion of posts expressing frame $f$ from 7 to 3 days before election $i$.
  • $\text{{post}}_{fi}$: mean proportion of posts expressing frame $f$ from 3 to 7 days after election $i$.
@@ Results
Aiming at defining a baseline distribution of frame usage across unions, we first looked at the distributions of posts by union. We identified skewness (minimum = 779, 25th percentile = 1,921, 75th percentile = 5,442, maximum = 10,452, full distribution in Figure E1) and thus implemented a bootstrap-like down-sampling procedure. Over-represented unions were down-sampled to the minimum number of posts in the distribution, and the procedure was repeated with five different random seeds. For each frame, we computed the median proportion of posts expressing that frame across seeds. Overall, the median baseline distribution of frame usage is as follows: diagnostic in 40.9% of posts, community in 39.4%, prognostic in 28.4%, engagement in 23.1%, and motivational in 15.5%.
@@
Figure 1: Relative usage of discourse frames by labor unions compared to the baseline median (percentage difference over baseline). Red color indicates a usage higher than the baseline, blue color indicates a usage lower than the baseline. The dendrogram shows the clustering of unions based on their framing patterns across the five frame types. Union acronyms are colored according to their organizational structure.
@@
Figure 3: Posterior mean log-odds estimates for selected variables, averaged across ten random seeds. Black dots indicate posterior means; grey lines show 90% Highest Density Intervals (HDIs) for each seed. Ratios denote the number of seeds (out of ten) with HDIs excluding zero. Variables shown meet this criterion in at least six seeds.
@@
Figure 4: Changes in frame usage before and after elections, comparing lost and won cases. Bold values in the cells refer to the difference between proportion of pattern presence for a given frame between losses and wins. Raw counts of union-case instances assigned to each cell are reported for losses (L) and wins (W). The $\star$ sign indicates a significant difference between losing and winning ($p<0.05$).
@@
While our study makes both methodological and substantive contributions, several limitations should be noted. First, our data is observational and correlational; consequently, observed associations between communication features and electoral success cannot be interpreted causally. In particular, unobserved organizational characteristics or strategic anticipation of election outcomes may jointly influence both communication practices and results. Second, the analysis is limited to Facebook—an important but partial element of unions’ communication strategies, which also involve other digital platforms and offline channels. Out of over 500 unions with available election outcomes, only 66 had Facebook accounts, and only 40 provided downloadable data, resulting in a relatively small and potentially skewed sample. Consequently, our findings may not generalize to all unions, particularly those for which Facebook is not a primary communication channel. Third, our analysis focuses on posts from the week immediately before and after each election. This window captures communication closely tied to electoral events and avoids noise from overlapping elections for a given union. However, it may miss broader or longer-term communication strategies, so results should be interpreted with this temporal limitation in mind. Fourth, we focused on national and international level union accounts to ensure sufficient activity, which may obscure dynamics in smaller or more localized organizations. Changes in data access policies could allow for analyses that can estimate the impact of such issues. Fifth, we analyzed five discourse frames identified in prior work on social movements and civil society organizations. This enhances comparability but may overlook additional frames or nuances. Finally, unions differ in size, resources, and political context, factors likely shaping communication in ways not fully accounted for here. We also acknowledge that systematic analyses of union communication could be misused by political actors, employers, or other organizations, underscoring the need for careful ethical reflection when applying these methods.
@@
The National Labor Relations Board (NLRB) election data has been widely used to study the dynamics of unionization in the United States. A central line of research has focused on identifying predictors of union success. Early work by Maranto and Fiorito (1987) highlights union size, internal democracy, strike activity, and centralized decision-making as key determinants of certification outcomes. Extending this perspective, Ferguson (2008) builds a sequential model of organizing, tracing the process from election petitions to first contracts, and identifies union size and unfair labor practice charges as critical factors across stages.
A second strand of studies combines NLRB data with external sources to investigate how broader institutional and demographic contexts shape unionization. For example, Sojourner and Yang (2022) links NLRB elections with Occupational Safety and Health Administration data to show how certification influences the enforcement of workplace safety laws. Similarly, Ferguson (2015) integrates Equal Employment Opportunity Commission surveys to examine how unionization affects workplace composition, with attention to racial and social diversity. Bronfenbrenner (2005) takes a related approach by comparing bargaining units across gender compositions, assessing the role of company characteristics, union demographics, and organizing tactics.
Other contributions place NLRB election outcomes within broader structural or political-economic dynamics. Stepan-Norris and Southworth (2010) analyzes competition among AFL-CIO affiliates, showing how inter-union rivalry shapes union density when accounting for political and economic conditions. Schaller (2023) instead focuses on long-term union decline, attributing it to sectoral and regional shifts in economic activity. Finally, Ferguson, Dudley, and Soule (2018) connects labor organizing to wider social movements, introducing the concept of osmotic mobilization to demonstrate how protest activity in society at large correlates with support for unions in representation elections.
This literature demonstrates the versatility of NLRB election data for analyzing both the internal dynamics of organizing and its intersection with external institutional, demographic, and political factors. While existing studies have focused on predictors of election success, workplace outcomes, and long-run trends, our approach incorporates the digital domain and employs this data as an anchor to investigate transformations in digital communication.
@@
With the large-scale adoption of digital technologies, labor unions have turned to online platforms as tools to engage with members, recruit new participants, and communicate with stakeholders. Scholars emphasize that unions’ strategies are shaped not only by external political contexts, but also by internal organizational choices (Martínez Lucio 2003). In this regard, Frege and Kelly (2004) identifies six strategic choices related to union revitalization, among which organizing—recruiting, representing, and retaining members, as well as expanding mobilization capacity—is particularly relevant in the digital context. Within these revitalization strategies, communication is central: discourse frames provide unions with a way to articulate grievances, propose solutions, and mobilize collective identities, thereby directly supporting organizing efforts. Building on this perspective, a consistent body of literature has linked framing and the labor movement.  Kelly (1998) adapts the collective action frames introduced by Snow, Benford et al. (1988) to the labor context and argues that successful organizing depends on workers interpreting grievances as unjust, identifying solutions, and believing collective action can be effective. Similarly, Voss and Sherman (2000) highlights the importance of identifying the right strategic repertoires of communication and engagement for the labor revitalization efforts to be successful. Other foundational work that followed highlights the importance of identifying the right frames for women and minorities (Bronfenbrenner 2005), and millenials (Milkman 2013).
Much of the literature on revitalization through digital media communication has focused on mobilization. Studies of the Fight for Fifteen movement, for example, use qualitative and mixed-methods analysis to demonstrate how digital media facilitate mobilizing efforts and illuminate the interplay between elements of collective action and the connective logic of grassroots initiatives (Pasquier and Wood 2018; Pasquier, Daudigeos, and Barros 2020; Frangi, Zhang, and Hebdon 2020). Similarly, Houghton and Hodder (2021) focuses on mobilizing frames across platforms, coding social media posts in terms of framing, attribution, and calls to action. Beyond mobilization, research has also examined how digital networks support organizational purposes, such as building collective voices prior to mobilization (Lazar, Ribak, and Davidson 2020), or reaching marginalized groups otherwise excluded from traditional organizing channels (Wood 2015; Thornthwaite, Balnave, and Barnes 2018). The effectiveness of communication strategies is also contingent on other important factors, such as organizational and leadership structures, as demonstrated in the comparative analysis of 2018 teachers’ strikes in Oklahoma and Arizona (Blanc 2022).
@@ Appendix C Topic Modeling on Selected Unions
((a)) IAFF (Craft)
((b)) SEIU (Industrial)

### v3

@@ Introduction
https://github.com/ariannap13/laborunions
@@
  • $\text{{pre}}_{fi}$: mean proportion of posts expressing frame $f$ from 7 to 3 days before election ii.
  • $\text{{post}}_{fi}$: mean proportion of posts expressing frame $f$ from 3 to 7 days after election ii.
@@ Results
Aiming at defining a baseline distribution of frame usage across unions, we first looked at the distributions of posts by union. We identified skewness (minimum = 779, 25th percentile = 1,921, 75th percentile = 5,442, maximum = 10,452, full distribution in Figure E1) and thus implemented a bootstrap-like down-sampling procedure. Over-represented unions were down-sampled to the minimum number of posts in the distribution, and the procedure was repeated with five different random seeds. For each frame, we computed the median proportion of posts expressing that frame across seeds. Overall, the median baseline distribution of frame usage is as follows: diagnostic in 40.7% of posts, community in 39.4%, prognostic in 28.3%, engagement in 23.2%, and motivational in 15.4%.
@@
Figure 1: Relative usage of discourse frames by labor unions (percentage difference from baseline). Red indicates above-baseline usage, blue below. The dendrogram clusters unions by framing patterns across the five frames, and acronyms are colored by organizational structure.
@@
Figure 3: Posterior mean log-odds for selected variables, averaged over ten seeds. Black dots: means; grey lines: 90% HDIs per seed. Ratios: number of seeds with HDIs excluding zero (shown if $\geq 6$ seeds).
@@
Figure 4: Changes in frame usage pre- vs. post-election, by lost and won cases. Bold values: difference in frame proportions between losses and wins. Counts: union-case instances (L=loss, W=win). $\star$: significant difference ($p<0.05$).
@@
While our study makes both methodological and substantive contributions, several limitations should be noted. First, our data is observational and correlational; consequently, observed associations between communication features and electoral success cannot be interpreted causally. In particular, framing may reflect underlying organizational capacity rather than independent strategic choice: unions with greater resources, stronger leadership, or more intensive offline organizing may both communicate differently and be more likely to win elections. Second, the analysis is limited to Facebook, an important but partial element of unions’ communication strategies, which also involve other digital platforms and offline channels. Out of over 500 unions with available election outcomes, only 66 had Facebook accounts, and only 40 provided downloadable data, resulting in a relatively small and potentially skewed sample. Consequently, our findings may not generalize to all unions, particularly those for which Facebook is not a primary communication channel. Third, our analysis focuses on posts from the week immediately before and after each election. While this window is intended to capture communication tied to electoral events, it may be too narrow to reflect longer-term framing strategies or too broad, capturing stable communication patterns rather than election-specific adjustments. To mitigate the latter concern, we account for longer-term trends in frame usage by leveraging periods without elections. Fourth, we focus on national and international level union accounts to ensure sufficient activity, which may obscure dynamics in smaller or more localized organizations. Changes in data access policies could allow for analyses that can estimate the impact of such issues. Fifth, we analyzed five discourse frames identified in prior work on social movements and civil society organizations. This enhances comparability but may overlook additional frames or nuances. Finally, unions differ in size, resources, and political context, factors likely shaping communication in ways not fully accounted for here. We also acknowledge that systematic analyses of union communication could be misused by political actors, employers, or other organizations, underscoring the need for careful ethical reflection when applying these methods.
@@
The National Labor Relations Board (NLRB) election data has been widely used to study unionization dynamics in the United States, particularly predictors of union success. Early work by Maranto and Fiorito (1987) highlights union size, internal democracy, strike activity, and centralized decision-making as key determinants of certification outcomes. Extending this perspective, Ferguson (2008) models organizing as a sequential process from petitions to first contracts, highlighting union size and unfair labor practice charges as critical across stages.
A second strand of studies combines NLRB data with external sources to investigate how broader institutional and demographic contexts shape unionization. For example, Sojourner and Yang (2022) connects certification to workplace safety enforcement, while Ferguson (2015) studies workplace composition using Equal Employment Opportunity Commission data. Bronfenbrenner (2005) takes a related approach by comparing bargaining units across gender compositions, assessing the role of company characteristics, union demographics, and organizing tactics.
Other studies situate elections within broader political-economic dynamics. Stepan-Norris and Southworth (2010) analyzes inter-union competition, while Schaller (2023) links union decline to sectoral and regional shifts. Finally, Ferguson, Dudley, and Soule (2018) connects labor organizing to wider social movements, introducing the concept of osmotic mobilization to demonstrate how protest activity in society at large correlates with support for unions in representation elections.
This literature demonstrates the versatility of NLRB election data for analyzing both the internal dynamics of organizing and its intersection with external institutional, demographic, and political factors. While prior work focuses on predictors of success, workplace outcomes, and long-term trends, we incorporate the digital domain, using NLRB data to anchor the analysis of online communication.
@@
With the widespread adoption of digital technologies, labor unions increasingly use online platforms to engage members, recruit participants, and communicate with stakeholders. Union strategies are shaped by both external political contexts and internal organizational choices (Martínez Lucio 2003). Among revitalization strategies, organizing—recruiting, representing, retaining members, and expanding mobilization capacity—is particularly prominent in digital settings (Frege and Kelly 2004). Communication plays a central role in these efforts, as discourse frames help articulate grievances, propose solutions, and mobilize collective identities.
A substantial body of work links framing to labor organizing. Building on collective action theory (Snow, Benford et al. 1988)Kelly (1998) argues that successful mobilization depends on workers perceiving injustice, identifying solutions, and believing in collective efficacy. Similarly, Voss and Sherman (2000) highlights the importance of identifying the right strategic repertoires of communication and engagement for the labor revitalization efforts to be successful. Other foundational work that followed highlights the importance of identifying the right frames for women and minorities (Bronfenbrenner 2005), and millenials (Milkman 2013).
Research on digital communication has largely focused on mobilization. Studies of the Fight for Fifteen movement show how social media supports mobilization and connects collective action with grassroots dynamics (Pasquier and Wood 2018; Pasquier, Daudigeos, and Barros 2020; Frangi, Zhang, and Hebdon 2020). Other work analyzes mobilizing frames across platforms (Houghton and Hodder 2021), coding social media posts in terms of framing, attribution, and calls to action. Beyond mobilization, research has also examined how digital networks support organizational purposes, such as building collective voices prior to mobilization (Lazar, Ribak, and Davidson 2020), or reaching marginalized groups otherwise excluded from traditional organizing channels (Wood 2015; Thornthwaite, Balnave, and Barnes 2018). The effectiveness of communication strategies is also contingent on other important factors, such as organizational and leadership structures, as demonstrated in the comparative analysis of 2018 teachers’ strikes in Oklahoma and Arizona (Blanc 2022).
@@
Acknowledgements
This work was supported by the Carlsberg Foundation through the COCOONS project (CF21-0432), and partially supported by the National Science Foundation (Grant IIS-1815875). Funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@ Appendix C Topic Modeling on Selected Unions
(a) IAFF (Craft)
(b) SEIU (Industrial)